### v1

# MoMentS: A Comprehensive Multimodal Benchmark for Theory of Mind

Abstract
Understanding Theory of Mind is essential for building socially intelligent multimodal agents capable of perceiving and interpreting human behavior. We introduce MoMentS (Multimodal Mental States), a comprehensive benchmark designed to assess the ToM capabilities of multimodal large language models (LLMs) through realistic, narrative-rich scenarios presented in short films. MoMentS includes over 2,344 multiple-choice questions spanning seven distinct ToM categories. The benchmark features long video context windows and realistic social interactions that provide deeper insight into characters’ mental states. While the visual modality generally enhances model performance, current systems still struggle to integrate it effectively, underscoring the need for further research into AI’s multimodal understanding of human behavior.
MoMentS: A Comprehensive Multimodal Benchmark for Theory of Mind
[…]
Figure 1: Overview of MoMentS questions.
1 Introduction
Throughout our lives, we continuously generate hypotheses about other people’s emotions, knowledge, and a range of other mental states; these hypotheses guide how we understand and interact with others. This ability, known as Theory of Mind (ToM) Premack and Woodruff (1978), is essential for interpreting behavior at the individual level and fundamental to coherent human social interaction Byom and Mutlu (2013).
Humans rely on more than just language to express their mental states. Gaze, facial expressions, body posture, gestures, and vocal cues all play an important role in communicating how we feel and what we think. This combination of verbal and non-verbal cues provides relevant multimodal information to infer mental states of others Byom and Mutlu (2013); Bayliss and Tipper (2006); De Sonneville et al. (2002).
For artificial agents, this information can serve as multimodal input that enhances socially intelligent behavior, empowering users across a wide range of applications: from facilitating communication and enhancing collaboration to offering companionship. A robust ToM enables such systems to anticipate intentions, understand desires and emotions, and detect knowledge gaps, to adapt their behavior to support users more effectively Oguntola et al. (2021). Importantly, this requires not only inferring individual mental states, but doing so in context—accurately "reading the room" by processing these signals to interpret human behavior in socially situated settings Williams et al. (2022).
Most existing benchmarks proposed to measure ToM in artificial agents predominantly center around belief-tracking tasks within text-based narratives or simplified multimodal settings Chen et al. (2025a). While these approaches evaluate models’ ability to reason about who knows or believes what, they frequently neglect the interplay of emotions, intentions, pragmatic communication, and social contexts that characterize genuine human interactions. Consequently, a clear gap exists between existing evaluations and the richer, socially grounded reasoning required in realistic scenarios.
To support the development of socially intelligent multimodal agents and assess current models’ ToM in realistic, socially grounded scenarios, we introduce MoMentS (Multimodal Mental States), a comprehensive multimodal video question-answering benchmark designed to evaluate ToM across seven abilities derived from the ATOMS taxonomy Beaudoin et al. (2020): Intentions, Desires, Beliefs, Knowledge, Percepts, Non-literal Communication, and Emotions. The dataset comprises 2,344 human-annotated questions and 9,376 candidate answers sourced from 168 long-form videos, annotated with short and long context windows, multimodal cue markers, and adversarially-generated distractors to minimize biases.
To the best of our knowledge, MoMentS is the first benchmark to evaluate multimodal ToM in real-world videos with human actors, framing it explicitly as a socially situated ability. Our contributions are as follows:
  • MoMentS: A novel multimodal benchmark with over 2,300 questions from real-world, long-form video data, explicitly structured to assess diverse ToM abilities.
  • An LLM-in-the-loop annotation framework designed to produce challenging distractors and mitigate bias in answer sets.
  • A baseline evaluation of multimodal LLMs, highlighting that although visual information improves performance, current models still predominantly rely on textual cues, underscoring the need for improved multimodal integration throughout the reasoning process.
2 Related Work
Prior benchmarks for ToM broadly fall into two categories: text-only and multimodal. Traditional text-only benchmarks, such as TOMI Le et al. (2019) and HI-TOM He et al. (2023), predominantly focus on probing models’ ability for nested belief tracking and logical inference through text stories lacking realistic social context. TOMBench Chen et al. (2024) expands beyond belief tracking alone, incorporating a broader taxonomy of social and pragmatic ToM tasks (e.g., faux-pas detection, persuasion, hidden emotions, desires) within everyday textual scenarios. Despite this richer coverage, it remains constrained by its purely textual format, lacking multimodal information critical to human social understanding Byom and Mutlu (2013).
Multimodal approaches such as MMToM-QA Jin et al. (2024) present procedurally-generated videos of single actors in household tasks, primarily evaluating goal and belief inferences without meaningful social interaction or emotional complexity. Similar to the text-only evaluations discussed above, this setup fails to reflect the depth and nuance of genuine human social behavior, limiting its applicability in evaluating socially intelligent AI systems Chen et al. (2025a).
From the social intelligence perspective, Social IQa Sap et al. (2019) probes social and emotional intelligence of models through multiple choice questions that require reasoning about social motivations, reactions, and actions based on specific situations. SOTOPIA Zhou et al. (2023) evaluates how models navigate complex social scenarios and achieve social goals. EmoBench Sabour et al. (2024) measures emotional intelligence by assessing models’ ability to understand and apply emotional knowledge in complex social scenarios. However, these works are again limited to text-only evaluations and do not measure ToM directly.
Social Genome Mathur et al. (2025) (based on SocialIQ2 Wilf et al. (2023)) addresses the evaluation of social interaction understanding in VLMs through video-based multiple-choice questions, but videos are limited to 60 second clips, and evaluation is not designed to evaluate ToM. Furthermore, Guo et al. (2023) observed a strong bias in the representations of correct and incorrect answer candidates, where LLMs can achieve high accuracy with no context or transcription.
Given the limitations in prior work, there is a need for evaluating ToM within realistic multimodal settings, capturing authentic social interactions beyond goals and beliefs alone Chen et al. (2025a).
3 Dataset Design
Recognizing the limitations of previous benchmarks, we design MoMentS based on two core principles: (1) an established taxonomy of socially relevant ToM abilities –Emotions, Non-Literal Communication, Desires, Intentions, Knowledge, Percepts, and Beliefs– to evaluate ToM beyond the commonly addressed belief and goal probing abilities, and (2) long-form videos with real human actors that provide sufficient context and multimodal signals (e.g., facial expressions, gaze, body language, speech tone) to characterize interpersonal dynamics and mental states. This section outlines our taxonomy for probing different ToM abilities, the video selection process, and the annotations included in each question.
Figure 2: MoMentS Annotation Pipeline. Different colored t-shirts represent different annotators/reviewers.
3.1 Taxonomy and Question Design
We adopt the ATOMS taxonomy (Abilities in Theory of Mind Space) introduced by Beaudoin et al. (2020) from their meta-analysis of ToM studies and proposed as a systematic framework for model evaluation by Ma et al. (2023). ATOMS categorizes ToM into seven distinct abilities: Knowledge, Emotions, Desires, Beliefs, Intentions, and Non-literal Communication (NLC). We describe each ability in Table 1. This taxonomy supports precise question formulation and provides a detailed framework for systematically evaluating specific ToM abilities in models. We design annotation guidelines (See Appendix A.9) around it.
ToM Ability Overview Example Q Example A
Knowledge Understanding what a person knows or does not know based on their sensory access. Why is the old man more interested in the box on the table than the book? The box is unfamiliar to him, therefore more interesting.
Emotions Identifying and reasoning about emotional responses, their evolution, and when emotions are hidden or complex. How is the girl feeling after hearing the news? She is saddened but hides her true feelings with a smile.
Desires Situations that involve preferences, conflicting desires, or actions driven by desire. Why is the man making an effort to defend his friend’s actions? He knows his friend stole, but feels attracted to him and wants to stay on good terms.
Beliefs Comprehending true and false beliefs and how beliefs influence actions. Why does the mother pretend to be surprised by the birthday cake? She thinks her daughter believes the cake is a surprise, even though she helped plan it.
Intentions Understanding goals, motivations, and the underlying reasons for actions. Why does the woman interrupt the conversation? She wants to change the topic.
Percepts Reasoning about what a character can or cannot perceive through their senses. Why does the old man not try to catch the falling vase? He did not see it, so he did not try to catch it.
NLC Interpreting humour, sarcasm, deception, and other speech that goes beyond literal meaning. Why does the man say he is stuck in traffic? He is lying to avoid conflict with his boss.
Table 1: Overview for ATOMS abilities covered in MoMentS with example question/answer pairs.
3.2 Video Selection
Existing datasets contain synthetic videos or minute-long clips to incorporate limited temporal context. We instead propose to use short films as these contain more complex characterizations and longer temporal contexts, while having a self-contained narrative. Our videos come from the SF20K dataset Ghermi et al. (2024), which contains a curated collection of short films from the YouTube channel Omeleto. Ghermi et al. (2024) verified that these films exhibit minimal information leakage to state-of-the-art language models compared to other common video sources like the sitcom Friends. Additionally, the videos are high-quality, vary in length (10 to 20 minutes), and provide complete, cohesive stories.
Not all short films have scenarios suitable for evaluating ToM. To filter these out, we prompted GPT-4o with film synopses to identify ones that likely contain interesting social dynamics. We then select videos with the highest likelihood of generating meaningful question-answer pairs and assign each annotator a a subset of these to annotate.
3.3 Data Organization
In line with prior work, we adopt a multiple-choice question-answer (MCQA) format, where each question includes one correct answer and three plausible but incorrect distractors. Figure 1 exemplifies two items from MoMentS, and more representative examples are presented in Appendix A.1. Below, we describe the structure and annotations included in each data point:
Questions
are derived from specific scenes in the short films and must probe one or more ToM abilities as defined in the ATOMS taxonomy.
Answer Set
includes one correct option and three distractors. Annotators are instructed to write distractors that are as plausible as possible, such that only a nuanced understanding of the context can reveal the correct answer. We paid special attention to the distractors, see Section 4.2 for more details on this.
Tags for ToM Abilities
specify which ToM abilities (See Table 1) are targeted by the question. Questions may be annotated with multiple abilities, acknowledging that these often intersect in various scenarios.
Timestamps
mark the start and end of the video segment relevant to the question. Each question is annotated with two context windows:
  • Full Context Window $[t_{0},t_{j}]$: A longer segment starting from the beginning of the video, intended to provide full narrative context useful for understanding character backgrounds, motivations, and evolving social dynamics.
  • Focused Context Window $[t_{i},t_{j}]$: A shorter segment containing only the immediate context required to answer the question. This window excludes broader narrative information, focusing instead on the specific scene being queried.
During evaluation, we explicitly instruct models that the question refers to the end of the provided interval ($t_{j}$). This approach minimizes reliance on temporal references that may hint at the correct answer, which requires understanding the interaction. If leveraged effectively, the Full Context Window provides all the information required to understand characters, providing better insights into their mental states and interpersonal dynamics.
Multimodal Cue Tags
indicate whether answering the question relies on interpreting specific nonverbal or auditory signals. These tags were optionally marked by annotators and are present only when such cues were deemed necessary for understanding the interaction. The possible cues include: "Facial Expressions or Gaze", "Body Language", and "Speech-related".
4 Annotation Methodology
(a) Boxplot of accuracies in the Focused Context Window.
(b) Boxplot of accuracies in the Full Context Window.
Figure 3: Boxplots comparing average accuracies across different models (at $\sim$7B parameter scale) and ToM abilities. Results for Audiovisual LLMs (▲) are reported using video and audio inputs ($VA$), while results for Video LLMs (⚫) use video and transcript inputs ($VT$).
Creating multiple-choice questions for this task is challenging. Annotators must understand different ToM abilities, find relevant moments in short films, and write clear questions. Making good distractors is also difficult because humans often create distractors that models can easily guess without seeing the video context, as observed by Guo et al. (2023) in other multimodal social understanding datasets.
To address these challenges, we conducted two pilot annotation rounds (see Appendix A.2) before launching the main annotation phase. Findings from the pilots helped us refine our pipeline to address the cognitive demands of ToM question creation, reduce annotation biases, and ensure question quality. The final methodology included carefully structured annotation phases, refined guidelines, and a custom-built platform to support robust distractor generation.
4.1 Annotation Pipeline
Statistic Length
Question Length $12.64\pm 4.2$
Correct Answer Length $14.62\pm 7.8$
Distractor Length $14.97\pm 7.7$
$[t_{i},t_{j}]$ length $42.44\pm 55.5$
$[t_{0},t_{j}]$ length $388.47\pm 262.3$
Number of Videos $168$
Video length (m) $14.56\pm 4.65$
Table 2: Top: Mean length $\pm$ SD of questions, correct answers, and distractors (in words), together with the average duration of the focused $[t_{i},t_{j}]$ and full $[t_{0},t_{j}]$ context windows (in seconds). Bottom: Number of Videos and average duration (in minutes)
Annotation guidelines were centered around the ATOMS taxonomy and the specific goals of the benchmark. They included illustrative examples, key indicators (what to look for) for each ToM ability, and clearly defined criteria for both acceptable and problematic question types. We iteratively refined the guidelines based on feedback from our expert sociologist and from the annotators themselves during the pilot runs.
The main annotation phase spanned six weeks and involved 16 annotators who collectively produced 2,344 questions. This phase followed the methodology developed during the second pilot and incorporated several design choices aimed at improving quality and reducing bias (see Figure 2 for an overview):
  • Annotators were asked to watch the full short film before writing questions to ensure understanding of character motivations and social dynamics.
  • Each was assigned 2–3 ToM abilities to specialize in, promoting category-specific expertise.
  • The schedule alternated weekly: a week focused on writing questions, the next on creating distractors for peers’ questions. During the distractor-creation stage, annotators flagged poorly written or overly subjective questions, adding a layer of peer-based quality control.
  • A custom platform integrated an LLM for real-time distractor feedback, flagging biased sets automatically (Section 4.2).
  • We provided weekly feedback based on a review of the submitted material. For questions, we emphasized clarity, appropriate ToM category assignment, and avoidance of overly subjective QA pairs. For distractors, we focused on ensuring that none of the distractors could be considered a "technically correct" answer.
  • We provided bonuses for early submissions and for the annotators who produced the highest-quality questions.
This approach encouraged focused annotation, peer-based quality control, and robust distractor generation, resulting in the final MoMentS evaluation dataset. Table 2 report statistics about the dataset. In Appendix A.6, we report the demographics of annotators, cost of the annotations, and number of questions associated to each ToM ability.
4.2 Framework for Distractor Creation
Models frequently rely on subtle biases to guess correctly; our initial pilot batch showed this issue with models consistently achieving non-trivial performance by identifying correct answers without the required context. Creating high-quality distractors remains challenging for annotators despite providing them with guidelines; even subsequent re-annotation of distractors by us similarly demonstrated persistent biases.
While various post-hoc strategies exist to mitigate distractor bias Ye and Kovashka (2021); Guo et al. (2023), we integrate bias prevention directly into the annotation workflow. We designed a custom annotation platform embedded with an LLM acting as an on-the-fly evaluator for newly proposed distractor sets.
Given a question with one correct answer and three proposed distractors, the platform evaluates potential biases distractor as described in Algorithm 1.
Input:
Question $Q$, correct answer $a^{*}$, distractors $D=\{d_{1},d_{2},d_{3}\}$, trials $N$, threshold $k$.
Output: Indicator of biased distractors
$c\leftarrow 0$;
for $i\leftarrow 1$ to $N$ do
       $A\leftarrow\text{shuffle}(\{a^{*}\}\cup D)$;
       $a\leftarrow\text{LLMAnswer}(Q,A)$;
       if $a=a^{*}$ then
             $c\leftarrow c+1$;
if $c\geq k$ then
       return flag biased;
Algorithm 1 Distractor Set Assessment
We establish empirically determined $k=5$ and $N=6$ to balance reliability and computational efficiency. A distractor set is flagged as biased if the model identifies the correct answer $k$ or more times out of $N$ trials. We initially employed GPT-4o-mini as the LLM for the first 800 questions; as we observed that the cost was relatively low, we decided to use GPT-4o for the remaining annotations.
5 Experimental Evaluations
We conduct experiments to evaluate the performance of current multimodal models in inferring mental states and to identify the factors that influence their performance. Specifically, we aim to answer: (i) How well do these models perform overall and across different ToM abilities? (ii) To what degree does visual information and context length impact performance? and (iii) How effective is our LLM-in-the-loop distractor creation platform at mitigating answer set biases? To this end, we report model accuracies on MoMentS, ablate the effect of the visual modality and context window length, and assess performance in a no-context setting against baselines lacking bias-mitigation mechanisms.
Model $[t_{0},t_{j}]$ $[t_{i},t_{j}]$
$T$ $VT$ $T$ $VT$
⚫ LLaVA-Video-7B 47.0 49.36 (+2.3) 45.6 52.01 (+6.4)
⚫ InternVL2.5 8B 46.0 46.63 (+0.6) 45.4 51.79 (+6.4)
⚫ LongVA-7B-DPO 41.0 44.24 (+3.3) 41.5 44.5 (+3)
⚫ Qwen2.5 VL 8B 41.3 38.05 (+-3.3) 38.4 44.33 (+5.9)
⚫ LLaVA-Video-72B 63.2 65.96 (+2.8) 62.1 67.66 (+5.6)
⚫ InternVL2.5 78B 53.3 61.09 (+7.8) 52.2 61.48 (+9.3)
$T$ $VT$ $T$ $VT$
▲ Qwen2.5-Omni-7B 46.8 53.41 (+6.6) 45.8 56.19 (+10.4)
▲ VideoLLaMA2-7B-AV 37.6 40.96 (+3.3) 38.4 43.13 (+4.7)
▲ MiniCPM-o 2.6 (8B) 47.8 ** 47.1 50.17 (+3.1)
$A$ $VA$ $A$ $VA$
▲ Qwen2.5-Omni-7B 44.41 52.69 (+8.3) 48.46 55.59 (+7.1)
▲ VideoLLaMA2-7B-AV 34.22 42.32 (+8.1) 34.22 43.6 (+9.4)
▲ MiniCPM-o 2.6 (8B) 39.5 ** 39.6 48.25 (+8.6)
$A$ $A$
▼ Kimi-Audio-7B 31.7 48.6
▼ Qwen2-Audio-7B 34.5 35.5
$VA$
Human 86
Table 3: Accuracy of different models on MoMentS, reported for both Full $[t_{0},t_{j}]$ and Focused $[t_{i},t_{j}]$ context windows. For Video LLMs (⚫), we show performance with transcripts only (T) and video+transcripts (VT). For Audiovisual LLMs (▲), we additionally report audio (A) and video+audio (VA) inputs. For Speech LLMs (▼), we report performance using audio only (A). *Human evaluation was conducted on a subset of 100 samples in the Full $[t_{0},t_{j}]$ setting using video and audio (VA) by a participant external to the annotation process.
5.1 Experimental Setup
We evaluate three types of multimodal LLMs: Video, Audiovisual, and Speech-based.
Video LLMs
These models process visual and text inputs. We evaluate LLaVA-Video 7B and 72B Zhang et al. (2024b), InternVL 2.5 8B and 78B Chen et al. (2025b), LongVA 7B Zhang et al. (2024a), and Qwen2.5 VL 7B Bai et al. (2025). Each model receives 64 uniformly sampled frames per question (see Appendix A.5 for frame count ablations).
Audiovisual LLMs
These models process visual, text, and audio inputs. We evaluate Qwen2.5-Omni 7B Xu et al. (2025), VideoLLaMa2 7B Cheng et al. (2024), and MiniCPM-o 2.6 (8B) Yao et al. (2024).
Speech LLMs
These models process audio and text inputs. We evaluate Kimi-Audio 7B KimiTeam et al. (2025) and Qwen2 Audio Chu et al. (2024).
All models are evaluated using the transformers library Wolf et al. (2020) with temperature set to 0. Under 70B paramenters models run on a single NVIDIA A100 GPU, whereas larger models (>70B) run on three NVIDIA H100 GPUs.
5.2 LLM Evaluation
Since Video LLMs process only vision and text, we transcribe dialogues using ASR through WhisperX using Whisper large-v2 as the backbone model (refer to Appendix A.4 for an evaluation on the performance of the ASR system).
We evaluate models under different input conditions: textual or audio dialogues only (Transcripts ($T$) or Audio ($A$)), and combined inputs (Video+Transcripts ($VT$) or Video+Audio ($VA$)). Additionally, we compare model performance across two temporal contexts: the Full Context Window $[t_{0},t_{j}]$ and the Focused Context Window $[t_{i},t_{j}]$.
Global Accuracy
Table 3 reports the global accuracy on MoMentS; we observe that video input improves performance in most cases. However, the gains are modest, indicating that current models may underutilize visual cues. Performance tends to drop when using the longer Full Context Window, we attribute this to the fact that long video understanding is still challenging for open models.
Accuracy by ToM Ability
Figure 3 presents boxplots with per-model scatter points, showing model accuracy across different ToM abilities under two context window conditions. As observed in the globlal accuracies, models perform better using shorter Focused Context Windows, though the impact of context length varies by ability. For example, accuracy is notably higher for Emotions and Intentions questions with shorter contexts, suggesting that these tasks rely more on immediate cues. Among results within the longer Full Context Window, Knowledge, Desires, and Non-literal Communication (NLC) questions perform relatively better, suggesting that longer context may be beneficial for understanding characters’ background and effectively answering these questions. For both of the context settings, Percepts and Beliefs remain the most challenging abilities. Future work should investigate how context window length affects human performance in this task. In Figure 4, we further analyze the role of visual input by averaging accuracies across Video LLMs. While the visual modality is not yet fully leveraged, it still contributes positively across all abilities, indicating its importance for ToM reasoning.
Figure 4: Average accuracies across Video LLMs highlighting the effect of the vision modality accuracy on different ToM abilities (Focused Context Window).
Multimodal Cues
$\Delta_{t_{j}-t_{0}}$ $\Delta_{VT-T}$ $\Delta_{VA-A}$
Body Langauge 3.71 7.92 8.05
F. Exp. and Gaze 2.01 6.5 6.79
Speech-related 1.77 3.62 6.32
Table 4: Effect of context length and visual input on questions marked as reliant on multimodal cues. $\Delta_{t_{j}-t_{0}}$ is the average accuracy difference between Focused and Full Contexts, with positive values indicating better performance in shorter intervals. $\Delta_{VT-T}$ and $\Delta_{VA-A}$ represent average accuracy gains from adding visual input (VT vs. T across Video and Audiovisual LLMs; VA vs. A across Audiovisual LLMs), both reported for the shorter context interval.
We further analyze performance on questions requiring multimodal understanding (facial expression or gaze, body language, and speech-related cues). As shown in Table 4, incorporating visual input and using a shorter context window generally improves performance, particularly for questions involving Body Language and Facial Expressions or Gaze. Speech-related questions benefit less from visual information when dialogues are provided as transcripts ($VT$), indicating stronger reliance on textual cues; however, vision seems to show a larger benefit in these questions when dialogues are provided directly as audio inputs in audiovisual models ($VA$).
5.3 Evaluation on Answer Set Bias
In this section, we evaluate the impact of using an LLM-in-the-loop design during the annotation pipeline, specifically for distractor creation. For MCQA-style ToM evaluation to be meaningful, questions should not be answerable without access to some form of context such as video, audio, or transcripts. However, as observed in our initial pilot and in prior work Guo et al. (2023), models often exploit biases in question-answer sets to guess the correct answer even without contextual input.
We assess the extent of this issue by comparing MoMentS to two baselines: our initial pilot (which did not use LLM assistance for distractor creation) and SocialIQ2, a similar video MCQA dataset. We prompt models with only the questions and answer options (without context) and measure their accuracy. As shown in Table 5, our proposed LLM-assisted distractor generation substantially reduces answer-set bias and lowers model accuracy by over 20 percentage points, highlighting the effectiveness of our approach.
By reducing biases in the answer sets, we create greater headroom for models to improve through reasoning rather than shortcut learning.
Model SIQ2-dev M-P1 MoMentS
Qwen2.5 VL 8B 52.49 60.59 36.05 (-24.54)
LongVA-7B-DPO 53.38 58.49 34.85 (-23.63)
LLaVA-Video-7B 56.2 59.48 40.10 (-19.38)
InternVL2.5 8B 51.43 55.39 36.26 (-19.13)
Table 5: Accuracy by guessing the correct answer, where models are not provided with any context about the question. M-P1 refers to our first pilot study, and SIQ2-dev to the development set of SocialIQ2 Wilf et al. (2023).
6 Open Challenges for Future Model Development
Our evaluations on MoMentS suggest that current limitations in multimodal ToM performance may stem not only from the reasoning capabilities of large language models, but also from how these systems access and process multimodal evidence. Our findings point to several technical factors that likely limit models’ ability to reason about mental states in socially rich scenarios. In this section, we outline four open challenges that, if addressed, could foster progress toward building better social multimodal agents.
Capturing Prosody and Ambient Sound in Audio
Transcripts alone omit environmental sounds and paralinguistic cues (speaker prosody, intonation), which support accurate inferences about Percepts, Emotions, Intentions, and Non-literal Communication. In addition, errors in the ASR propagate downstream. While some audio-native models like Kimi-Audio and Qwen2.5-Omni (audio-only setting, $A$) demonstrate slight advantages over transcripts when vision is not provided, most transcript-only models perform comparably or better when visual input is present (see Table 3). Future research should focus either on integrating rich audio descriptors into existing video-text pipelines or improving the current audio-processing capabilities of audiovisual models, especially for understanding longer conversational contexts.
Precise Vision–Speech Alignment
Answering Who said what, when? requires time-synchronised links between each utterance, the speaking character, and the surrounding visual context. Without such alignment, models cannot track which speakers possess which knowledge, nor can they exploit gaze, facial expressions, or body language that modulate dialogue meaning. The small gains we observe from adding vision (Table 3), and the limited improvements on questions marked as reliant on visual cues (Table 4), indicate that existing pipelines underutilize this channel.
Human-Centered Frame Selection
Uniform frame sampling risks missing short yet meaningful signals while wasting computation on redundant content. Simply increasing the frame rate is expensive and, as our ablation in Appendix A.5 shows, does not improve performance. Specialized frame sampling strategies that prioritise human-salient events (faces, hands, gaze shifts) are needed to capture the cues that observers actually rely on.
Structured Reasoning over Multimodal Evidence
Reasoning improves a wide range of text-only tasks, including ToM benchmarks. However, as Mathur et al. (2025) reports, asking VLMs to reason neither boosts accuracy nor yields human-aligned explanations for social MCQA in videos. We argue that effective multimodal reasoning may be bottlenecked by the three challenges above: inadequate audio representations, weak vision–speech alignment, and sub-optimal frame selection. Until models receive richer, better-organized evidence, additional reasoning steps are unlikely to help.
7 Conclusion
We introduced MoMentS, a benchmark that probes seven ToM abilities in realistic, long-form videos. It contains over 2,300 human-annotated MCQA items with substantially reduced biases in answer sets compared to prior datasets. From baseline experiments with Video, Audiovisual, and Speech LLMs we observe: (i) visual input offers consistent yet modest gains, indicating underutilization of visual cues; (ii) using audio inputs does not yield a noticeable improvement over transcripts, suggesting that current models still have struggle to effectively integrate this modality; and (iii) performance tends to drop on extended context windows, highlighting limitations in long-range video reasoning.
Based on these results, we identify several open challenges that likely constrain progress on multimodal ToM tasks, ranging from multimodal alignment and audio processing to frame selection and reasoning over multimodal evidence. Addressing these issues will be essential for developing AI systems capable of truly understanding, predicting, and responding to human mental states in real-world social settings.
Limitations
We adopt a multiple-choice QA format in MoMentS to streamline annotation and ensure consistent evaluation. While this design supports scalable benchmarking, it limits analysis of lower-level behavioral cues such as turn-taking, speech acts, or gesture dynamics as we do not provide annotations on them. Investigating the relation between these cues and specific ToM abilities remains an important direction for future work. Additionally, MoMentS uses static video data, which does not capture model performance in interactive or dynamic social environments. Extending evaluation to such settings is a promising but currently challenging task, as it would require reliably simulating complex, multimodal human behaviors. Finally, although using multiple annotators per question could reduce subjectivity, resource constraints limited us to one annotator per question. To mitigate this, we incorporated peer-checking during distractor creation and conducted multiple rounds of author review to ensure data quality and consistency.
Appendix A Appendix
A.1 Samples from MoMentS Across ToM Abilities
Figure 5 presents representative samples of MoMentS questions covering different ToM abilities. Each example includes the question, the full answer set (one correct option and three distractors), the targeted ToM abilities, and any multimodal cues identified by annotators as relevant for answering the question.
Figure 5: Samples from MoMentS Representing Each ToM Abilities.
A.2 Pilot Annotations
We conducted two pilot annotation phases prior to the main annotation batch to identify challenges and refine our pipeline.
First Pilot Annotation
We recruited annotators through Prolific, selecting participants who were native English speakers with a university degree. Each annotator was asked to create both questions and distractors covering all seven ToM abilities. This pilot produced 268 question–answer sets. From analyzing submissions from this annotation batch, we identified the following issues:
  • Many questions were low quality, some had grammatical issues, others focused on plot rather than ToM.
  • Models achieved over 50% accuracy without context, pointing to biases in the distractor sets (see Table 5).
  • Annotators often mislabeled the ToM ability, indicating limited understanding of the categories.
We traced these problems to the following causes:
  • Time constraints imposed by Prolific’s system created pressure that negatively impacted annotation quality.
  • Prolific communication channels made direct communication with annotators difficult, as they did not communicate their questions effectively.
  • Tasking annotators with all seven categories was overwhelming, leading to overall misclassification.
  • Most effort was spent on writing questions, resulting in weaker distractors.
  • Models could exploit biases in seemingly good distractors, without needing any context to answer.
Second Pilot Annotation
To address these issues, we made the following changes:
  • We directly hired seven undergraduate students from psychology and social sciences and used group messaging for better communication.
  • Each annotator was assigned only 2–3 ToM abilities to help them specialize.
  • Annotation was split into two phases: creating questions in the first week and distractors in the second. This was done to help annotators concentrate their efforts on writing high-quality questions first, then shift their focus to creating high-quality distractors.
  • A custom annotation platform with an LLM was introduced to automatically flag biased distractors (see Section 4.2).
  • Annotators were encouraged to spread their work throughout the week to reduce low-quality submissions due to pressure in last-minute submissions.
  • We provided weekly reviews and feedback to improve consistency and quality.
This second pilot resulted in 350 high-quality questions. Most of the design choices from this phase were carried over to the main annotation batch.
A.3 Prompt For Video Filtering
You are a film critic and psychologist with expertise in Theory of Mind (ToM) as described by the ATOMS taxonomy. Your task is to analyze the movie synopsis and captions below to determine how likely it is that the movie includes themes or questions related to Theory of Mind.
Theory of Mind involves understanding and attributing mental states to oneself and others. Consider the following key components:
1. Knowledge: Recognizing that characters hold organized information and mental representations that shape their understanding.
2. Emotions: Identifying complex emotional responses, including mixed or evolving emotions.
3. Desires: Understanding that characters may have varied and sometimes conflicting desires driving their actions.
4. Beliefs: Discerning true versus false beliefs and recognizing higher-order beliefs (beliefs about others beliefs).
5. Intentions: Inferring characters goals and the reasoning behind their actions.
6. Percepts: Noting how characters perceive their world differently based on their sensory experiences.
7. Non-literal Communication: Interpreting subtleties such as sarcasm, humor, or metaphors that imply meanings beyond the literal words.
Using this framework, please analyze the following content:
Movie Synopsis: {synopsis}
Movie Captions: {caption}
Based on your analysis, provide a probability (as an integer percentage between 0 and 100) indicating how likely it is that this movie involves Theory of Mind related questions or themes. Your answer should be only the integer value with no additional commentary. choose the best number that seems appropriate based on the data.
A.4 Evaluation on ASR quality
In this subsection, we describe our audio processing pipeline, present, and report its ASR performance on a subset of human-annotated transcripts.
ASR Pipeline
We use WhisperX Bain et al. (2023) to transcribe the short films. Its multilingual capabilities make it suitable for both English and non-English videos in our dataset. For speaker diarization, we employ PyAnnote Bredin (2023).
ASR Quality Evaluation
We evaluate the ASR pipeline using different base Whisper models on a subset of 50 human-transcribed videos, reporting global Word-Error Rate (WER) and Diarization Error Rate (DER. For global WER we concatenate each file’s reference and ASR transcripts lower-casing and punctuation removal and computing $\text{WER}=(S+I+D)/N$, where $S$, $I$, and $D$ are the numbers of substituted, inserted, and deleted words, and $N$ is the total number of reference words. For DER, we evaluate only within spans where the reference marks speech. The score is $\text{DER}=(T_{\text{missed}}+T_{\text{confusion}})/T_{\text{ref}}$, where $T_{\text{missed}}$ is reference speech with no ASR coverage, $T_{\text{confusion}}$ is overlapped speech attributed to the wrong mapped speaker, and $T_{\text{ref}}$ is the total duration of speech in the reference annotation. We report these in Table 6, while Whisper large-v3 scores the lowest average global WER, in practice we notice that it failed to transcribe some of the videos. This does not happen with large-v2, whose DER is the lowest; because of this, we opted for the latter as the chosen model for transcribing audio for the Video LLMs.
global-WER DER
base 0.362 0.412
large-v2 0.206 0.363
large-v3 0.166 0.409
Table 6: Comparison of average WER and DER across the three evaluated models.
A.5 Ablation on number of frames
Increasing the number of video frames increases computational cost, as most Video LLMs embed frame patches significantly extending the context length processed by the language model. To assess the tradeoff between context length and performance, we evaluate three models on 1,500 randomly selected MoMentS entries using 64 and 96 frames.
As shown in Table 7, increasing the number of frames does not lead to consistent improvements. In several cases, performance actually drops, likely due to redundancy or context saturation. Based on these results, we use 64 frames for all main evaluations in the paper.
$[t_{0},t_{j}]$ $[t_{i},t_{j}]$
Model $T$ $VT$-64 $VT$-96 $T$ $VT$-64 $VT$-96
LLaVA-Video-7B 46.33 47.7 (+1.4) 47.3 (+1.0) 44.45 50.7 (+6.3) 49.7 (+5.3)
LongVA-7B-DPO 40.94 45.5 (+4.5) 42.6 (+1.6) 41.19 42.9 (+1.8) 44.6 (+3.4)
InternVL2.5 8B 45.58 45.6 (+0.1) 48.2 (+2.6) 44.45 51.7 (+7.3) 50.4 (+6.0)
Table 7: Global accuracy on a subset of 1,500 MoMentS samples using only transcripts ($T$), and transcripts plus 64 or 96 frames ($VT$-64 and $VT$-96). Results are reported for both the Full ($[t_{0},t_{j}]$) and Focused ($[t_{i},t_{j}]$) Context Windows. We mark in bold the highest increase over $T$ between 96 and 64 frames.
A.6 Dataset Statistics and Annotation Cost
The main annotation batch involved 16 participants: 12 undergraduate students in psychology and social sciences, two computer scientists, and two clinical psychologists. 12 of them were female and 4 male, all of them between 20 and 30 years old. Twelve of the annotators were from Canada, and the remaining were from Mexico. All participants were explained the purpose of their annotations in an onboarding session.
Annotation Cost
Annotators were compensated at a rate of 17 CAD per hour through UpWork. To encourage steady progress, a weekly bonus of 10 USD was provided to those who completed at least half of their assignments by midweek. An additional performance-based bonus of 150 USD was awarded to annotators who produced the highest-quality annotations. The total cost of the MoMentS main annotation effort amounted to 8,745 USD.
Dataset Statistics
ToM Ability # Questions
Emotions 599
Beliefs 379
Desires 386
Intentions 1026
Percepts 316
Knowledge 329
NLC 222
Table 8: Number of questions associated with each ToM ability.
Language Number of Videos
English 144
Russian 6
Spanish 5
French 3
Persian 3
Italian 1
Arabic 1
Swedish 1
Korean 1
Danish 1
Hindi 1
Japanese 1
Total 168
Table 9: Number of videos per language.
MoMentS contains 2,344 questions across 168 short films, the majority of which are in English (144). We also include a subset of 11 films in other languages. Table 9 reports the number of videos per language.
In Table 2, we report the average question length, average answer lengths, and durations of the full and focused context windows. We also display the distributions of lengths for answers, focused, and full context windows in Figures 6, Figure 7, and Figure 8, respectively.
A.7 Copyright and License
We release MoMentS annotations under a CC BY-NC-SA 4.0 license (Attribution-NonCommercial-ShareAlike 4.0 International), intended only for academic research purposes.
Following Ghermi et al. (2025) and Wilf et al. (2023), we do not distribute the video content directly. We provide URLs linking to the original videos on YouTube, complying with YouTube’s Terms of Service (https://www.youtube.com/static?template=terms).
A.8 Ethical Considerations
Representation and Bias
Most of MOMENTS videos are in English and reflect Western cultural norms. Additionally, annotators were from Canada and Mexico, which may influence interpretations of emotions, intentions, or non-literal communication.
Potential Misuse
MOMENTS is designed to evaluate models’ ability to infer mental states in socially grounded scenarios to foster progress in socially intelligent AI. However, ToM capabilities could also be misused to simulate deceptive, manipulative, or persuasive behavior in artificial agents. To mitigate this risk, we license the dataset for academic research only under a CC BY-NC-SA 4.0 license, and we strictly stand against any use in applications that exploit it for unethical purposes.
Personally Identifying Information or Offensive Content
Questions and answer sets do not contain personally identifying information as they use descriptors to refer to the characters. Since questions ask about character’s mental states, they do not contain offensive content.
Figure 6: Distribution of lengths for correct answers and distractors.
Figure 7: Length distribution of the Focused Context Windows.
Figure 8: Length distribution of the Full Context Windows.
A.9 Guidelines for Question and Distractor Annotation
The following pages contain the annotation guidelines provided to annotators during the first annotation batch. Separate documents were provided for the question creation and distractor creation stages to reflect the specific goals and challenges of each.
[Uncaptioned image] [Uncaptioned image]